what did you do to get the next three numbers after these 1,1,2,3,5,8,

Answer by ankor@dixie-net.com(22740)

get the next three numbers after these 1,1,2,3,5,8,
:
This looks like the Fibonacci series.
The next number is the sum of the two previous numbers, therefore:
the next number will be 5+8 = 13,
then 13+8 = 21
and 21+13 = 34
:
The first 14 numbers in the series would be:
1,1,2,3,5,8,13,21,34,55,89,144,233,377